OpenStudy (anonymous):

Write the expression as a single natural logarithm. 3ln3 + 3lnc A. ln(27 + c3) B. ln9c3 C. ln27c D. ln27c3

zepdrix (zepdrix):

Use this rule of logarithms to deal with the 3's in front. \[\huge a*\log(b)=\log(b^a)\] Then use this rule to combine your logs. \[\huge \log(b)+\log(c)=\log(b*c)\]

OpenStudy (anonymous):

@zepdrix i believe its B am i correct

zepdrix (zepdrix):

Mmmmm no.

zepdrix (zepdrix):

After you apply the first rule of logs, you should end up with a 3^3 term. Which is 27.